OpenStudy (anonymous):

Help! A swimming pool can hold a max of 360 gallons of water. The full pool develops a leak and is losing water at a constant rate. After 3 hours, the pool has 354 gallons of water in it. Write a function for g(t), the total number of gallons of water in the pool that is written in terms of t, the amount of time measured in hours that the leak developed.

OpenStudy (anonymous):

@timo86m hi tim i see youre online, any chance you could help me with this problem?

OpenStudy (anonymous):

oh that was easier than i thought. and if the leak is fixed after 20 hours and the owner immediately begins to refill the pool at a rate of 4 gallons of water per hour, write a piecewise functions for p(t), the total # of gallons of water that is in the pool for any time t, measured in hours.

OpenStudy (anonymous):

360-g(20) + 4 t when t>20

OpenStudy (anonymous):

wait is 360-g(20) + 4 t when t>20 the answer or is it the first one you posted?

OpenStudy (anonymous):

the second one is right :)

OpenStudy (anonymous):

oh ok and sorry one more question, if we solve for g(20), what does that represent?

OpenStudy (anonymous):

use the formula i gave you g(t)=118*t g(20)=118*20

OpenStudy (anonymous):

yes and i get 2360, but what does that represent?

OpenStudy (anonymous):

oh wait i did it wrong read it wrong lol

OpenStudy (anonymous):

haha darn

OpenStudy (anonymous):

it is g(t)=3/2 *t

OpenStudy (anonymous):

that is the function and g(t)=118t is wrong?

OpenStudy (anonymous):

yes i thought it meant after 3 hours it lost 354 not it went fro m360 to 354 i adjusted for it now tho :)

OpenStudy (anonymous):

oh ok so how did you get 3/2?

OpenStudy (anonymous):

so now your piece wise function is 360-g(t) when t is 0 to 20 360-g(20)+4*t when t>20

OpenStudy (anonymous):

ok and where did you get 3/2 from? and what does g(20), which is 30, represent?

OpenStudy (anonymous):

3/2 comes from (360-354)/3=3/2

OpenStudy (anonymous):

30 is the amount of water leak after 20 hours

OpenStudy (anonymous):

ohh ok but isnt 360-354=6 and 6/3=2?

OpenStudy (anonymous):

yes good catch lol i typed it in as 6/4 in my calculator

OpenStudy (anonymous):

btw your piece wise funtion is still right :)

OpenStudy (anonymous):

so now your piece wise function is 360-g(t) when t is 0 to 20 360-g(20)+4*t when t>20

OpenStudy (anonymous):

haha ok cool does and lol i was just going to ask that!

OpenStudy (anonymous):

because g(t) can later be modified and adjusted :)

OpenStudy (anonymous):

alright i got it thank you very much! :)

OpenStudy (anonymous):

do you think you could help me with some others if i post it?

OpenStudy (anonymous):

i'll be way slow but i can try lol.

OpenStudy (anonymous):

haha its ok! and thanks! ill post it in another question

OpenStudy (anonymous):

ok i posted it